# Analysis and Optimized CXL-Attached Memory Allocation for Long-Context LLM Fine-Tuning

### v1

@@ V-A1 Hardware and Software Specification
Experiments run on a server whose hardware and software stack are summarized in Table II. The platform combines a high‑performance CPU, ample local DRAM, and two cutting‑edge GPUs optimized for LLM workloads. CXL memory expansion is evaluated in two configurations: a single‑AIC setup and a dual‑AIC setup designed to reveal scalability limits and bandwidth‑contention effects. Both add‑in cards were SMART Modular devices [30]. Notably, even though the current configuration is limited to the availability of CXL AICs, the observations made in this study are representative and transferable to other CXL configurations.
@@
V-A2 Workload Setup
Data placement across local DRAM and CXL‑attached memory is managed by libnuma, which is exposed to PyTorch through a custom extension. DeepSpeed [45], which is the implementation of ZeRO-Offload, handles CPU offloading, while Flash‑Attention, Liger‑Kernel, and activation checkpointing enable efficient long‑context processing. Checkpointed activations, once generated, are offloaded to host DRAM. The study focuses on fine‑tuning large language models under a Causal Language Modeling objective. Two representative models—Qwen2.5‑7B [46] and Mistral NeMo 12B [47]—serve as workloads for exploring performance and scalability across varying context lengths, batch sizes, GPU counts, and AIC configurations. All experiments employ bf16 mixed‑precision training. The Adam optimizer maintains fp32 master parameters and optimizer states on the CPU, delivering the numerical stability required for LLM fine‑tuning. Details on specific context lengths, batch sizes, and AIC setups appear in the individual results subsections.
@@ V-B Performance Evaluation with a Single CXL AIC
(b) 12B model in a single-GPU scenario
(c) 7B and 12B models in a dual-GPU scenario
Figure 9: Training throughput comparison for three configurations: (1) Baseline (using local DRAM only), (2) Naive CXL adoption for single AIC, and (3) Using CXL-attached memory with CXL-aware Memory Allocation for single AIC. Performance is shown across varying context lengths and batch sizes for each scenario.
The single‑AIC scenario, corresponding to Config. A in Table II, serves as the starting point. Three configurations quantify the impact of adopting CXL‑attached memory: (1) Baseline, where all data reside in local DRAM; (2) Naive CXL, which combines 128 GiB of local DRAM with 512 GiB of CXL memory under a naive numactl interleave‑all policy; and (3) Our CXL, which pairs the same capacities while applying the proposed CXL‑aware Memory Allocation strategy. Figure 9(a) reports single‑GPU throughput for a 7B model across context lengths from 4K to 32K and batch sizes from 1 to 32. Relative to the baseline (normalized to 100%), the naive CXL setup drops to roughly 76%–94%. Degradation varies with workload mix: scenarios dominated by forward (FWD) and backward (BWD) passes suffer less because the latency‑critical optimizer step (STEP) forms a smaller share of runtime. Applying CXL‑aware allocation restores throughput to about 97%–99%, eliminating most of the loss. In other words, the corresponding performance drop is only 1% - 3% compared to the DRAM-only baseline and improved from the naive CXL by up to 21%.
Figure 9(b) shows analogous results for a 12 B model. The naive CXL configuration falls to roughly 72%–93% of the baseline, whereas the CXL‑aware method climbs to approximately 88%–96%. Even with the larger model pressing local memory to its limits, the drop remains confined to 4%–12%, a clear improvement over the naive policy. Dual‑GPU results appear in Figure 9(c) for both 7B and 12B models. Performance under the naive CXL policy reaches only 84%–94% of the baseline, while CXL‑aware allocation raises that range to 86%–99%. Although the gains are smaller than in single‑GPU trials—bandwidth contention still exists when two GPUs share a single AIC—the drop is limited to 1%–14%, depending on model size and context length. Multi‑AIC Striping, discussed next, addresses this remaining bottleneck by distributing traffic across multiple cards.
V-C Performance Evaluation with a Dual CXL AIC
(a) 12B model in a single-GPU scenario
@@ V-C Performance Evaluation with a Dual CXL AIC
(c) 12B models in a dual-GPU scenario
Figure 10: Training throughput comparison for three configurations: (1) Baseline (using local DRAM only), (2) Naive CXL adoption for dual AICs, and (3) Using CXL-attached memory with CXL-aware Memory Allocation and Multi-AIC Striping for dual AICs. Performance is shown across varying context lengths and batch sizes for each scenario.
The evaluation next turns to the dual‑AIC scenario, designated Config. B in Table II. Three configurations establish the impact of CXL memory: (1) Baseline, where all data remain in local DRAM; (2) Naive CXL, which pairs 128 GiB of DRAM with two 256 GiB AICs under a naive numactl interleave‑all policy; and (3) Our CXL, which uses the same capacities but applies CXL‑aware Memory Allocation plus Multi‑AIC Striping.
Figure 10(a) presents single‑GPU throughput for a 12B model across the same range of context lengths and batch sizes used earlier. Naive CXL drastically reduces throughput, delivering a reduction 2% - 11% of baseline performance. In contrast, CXL-aware allocation combined with Multi-AIC Striping restores the speed to 100% - 101% of the baseline, proving that dual cards can fully recover and slightly exceed native DRAM performance when bandwidth is balanced correctly. Figures 10(b) and 10(c) extend the study to dual‑GPU training of 7 B and 12 B models. Under the naive policy, throughput again reduces 2%–9% of the baseline. Introducing CXL‑aware allocation together with striping trims performance by at most 1%, effectively matching the DRAM‑only baseline even while two GPUs share system memory.
Earlier single‑AIC experiments showed limited gains because alone card became a bandwidth bottleneck. Dual‑AIC striping eliminates this constraint by distributing traffic, allowing CPU and GPU workloads to exploit the aggregate bandwidth of both cards. The resulting speedups over both single‑AIC tests and naive dual‑AIC adoption underline the importance of coordinated data placement and balanced striping when scaling CXL‑based memory expansion.
@@ VI Related Works
Offloading strategies have emerged as an effective way to break the GPU memory ceiling, enabling the training of models that would otherwise exceed on‑device capacity by staging tensors in CPU DRAM or NVMe SSDs [20, 21, 22, 23, 24]. Among these, the ZeRO series [20, 21] has become the most widely adopted, thanks to continuous maintenance that fixes bugs, adds support for new models, and preserves interoperability with complementary optimizations such as Flash‑Attention [40] and Liger‑Kernel [42]. Its reference implementation now underpins popular training frameworks, including Accelerate [48] and MS‑Swift [49]. Complementary research broadens the design space of memory‑centric training. Huan et al. [22] tailor an offloading mechanism specifically for large‑scale language models, while Chen et al. [23] evaluate pragmatic orchestration policies that coordinate CPU and GPU memory during fine‑tuning. Zeng et al. [24] further extend the idea to fully heterogeneous environments, automatically balancing both compute and memory resources.
For CXL-attached memory, several tiered‑memory systems (TMS) tackle the same capacity challenge at the operating system level. Maruf et al.’s TPP [34] classifies pages as hot or cold and places them in the most appropriate tier, keeping latency‑critical data in DRAM and relegating infrequently accessed data to CXL‑attached memory. Xiang et al.’s NOMAD [50] augments Linux with transactional page migration and page shadowing, permitting a page to be present in fast and slow memory simultaneously and moving it asynchronously outside the application’s critical path. Workload‑aware refinements continue to emerge. Vuppalapati et al. [51] observe that DRAM latency can spike near bandwidth saturation, sometimes exceeding that of CXL memory; their Colloid policy dynamically balances traffic across tiers to equalize effective access latency. San et al.’s M5 [52] embeds FPGA‑based hot‑page and hot‑word trackers in the CXL controller, supplying fine‑grained access statistics with negligible overhead. General‑purpose TMS designs operate transparently and require no application changes, yet this very neutrality can leave performance on the table for specialized workloads.
@@
To address the challenges of high latency and limited bandwidth in CXL-attached memory compared to local DRAM during long-context LLM fine-tuning, particularly the performance degradation in CPU-intensive computations and the bandwidth contention among multiple GPUs, this study is among the first to propose two key approaches. First, CXL-aware memory allocation strategically places latency-sensitive data in local DRAM and latency-tolerant data in CXL-attached memory based on access patterns. Second, multi-AIC striping distributes data across multiple CXL AICs to aggregate bandwidth, benefiting CPU-intensive phases and multi-GPU configurations. These optimizations significantly mitigate the performance drawbacks of CXL-attached memory, consistently outperforming naive CXL adoption with improvements ranging up to 21% in various scenarios. In particular, in the dual AIC configuration, our approach retains near-baseline performance, reducing the performance gap between DRAM-only setups to just 1% at most. These results demonstrate that this study enables CXL-attached memory as a viable solution for long-context LLM fine-tuning.

### v2

@@ V-A1 Hardware and Software Specification
Experiments run on a server whose hardware and software stack are summarized in Table II. The platform combines a high‑performance CPU (e.g., Intel Xeon 6780E), ample local DRAM (e.g., 512 GB DDR5), and two cutting‑edge GPUs (e.g., NVIDIA H100) optimized for LLM workloads. CXL memory expansion is evaluated in two configurations: a single‑AIC setup and a dual‑AIC setup designed to reveal scalability limits and bandwidth‑contention effects. Both add‑in cards were SMART Modular CMM-CXL-2.0 devices [25]. Notably, even though the current configuration is limited to the availability of CXL AICs, we argue that the observations made in this study are representative and transferable, as they are governed by the fundamental bandwidth and latency characteristics of the CXL protocol itself.
V-A2 Workload Setup
Data placement across local DRAM and CXL‑attached memory is managed by libnuma, which is exposed to PyTorch through a lightweight custom extension that intercepts memory allocation calls to enforce NUMA policies. DeepSpeed [38], which is the implementation of ZeRO-Offload, handles CPU offloading, while Flash‑Attention, Liger‑Kernel, and activation checkpointing enable efficient long‑context processing. Checkpointed activations, once generated, are offloaded to host DRAM. The study focuses on fine‑tuning large language models under a Causal Language Modeling objective. Two representative models, including Qwen2.5‑7B [39] and Mistral NeMo 12B [40], serve as workloads for exploring performance and scalability across varying context lengths, batch sizes, GPU counts, and AIC configurations. All experiments employ bf16 mixed‑precision training. The Adam optimizer maintains fp32 master parameters and optimizer states on the CPU, delivering the numerical stability required for LLM fine‑tuning. Details on specific context lengths, batch sizes, and AIC setups appear in the individual results sections.
@@ V-A2 Workload Setup
(a) 7B model in a single-GPU scenario
(b) 7B model in a dual-GPU scenario
(c) 12B models in a single-GPU scenario
(d) 12B models in a dual-GPU scenario
Figure 14: Training throughput comparison of three dual-AIC configurations: Baseline (local DRAM only), Naive CXL (interleaving), and CXL-Aware Allocation (labeled as Our), evaluated across varying models, context lengths and batch sizes.
@@ V-B Performance Evaluation with a Single CXL AIC
The single-AIC setup, corresponding to Config. A in Table II, serves as the baseline environment. Three configurations are evaluated to quantify the impact of CXL-attached memory. The first is the baseline, where all data reside in local DRAM. The second, naive CXL, combines 128 GiB of local DRAM with 512 GiB of CXL memory using a uniform numactl --interleave=all policy. The third, CXL-aware allocation, employs the same capacities but applies the proposed fine-grained memory allocation extension and CXL-aware memory allocator.
Figure 13(a) shows single-GPU throughput for a 7 B model under varying context lengths (4 K–32 K) and batch sizes (1–32). Relative to the baseline (normalized to 100%), the naive CXL configuration sustains only 76%–94% throughput, depending on workload mix. Workloads dominated by forward (FWD) and backward (BWD) passes experience smaller losses because the latency-sensitive optimizer (STEP) occupies a smaller runtime fraction. By contrast, CXL-aware allocation restores throughput to 97%–99%, reducing the degradation to just 1%–3% compared to DRAM-only and outperforming the naive policy by up to 21%.
Figure 13(b) presents dual-GPU results for the same 7 B model. The naive CXL setup achieves 77%–93% of baseline throughput, while CXL-aware allocation improves this to 93%–97%, narrowing the gap to 3%–7%. The smaller gain compared with the single-GPU case stems from bandwidth contention: both GPUs share a single AIC, effectively halving available read/write bandwidth. This bandwidth constraint limits both FWD and BWD phases, underscoring the benefit of deploying multiple AICs. Even with identical total capacity, multiple AICs aggregate bandwidth more effectively, allowing each GPU to access sufficient bandwidth and thus recover full performance. Section V-C further analyzes this multi-AIC configuration.
Figure 13(c) extends the analysis to a 12 B model under a single-GPU setup. The naive CXL configuration reaches 72%–93% of baseline throughput, while CXL-aware allocation raises this to 88%–96%, reflecting a 4%–12% shortfall relative to DRAM-only but up to 16% improvement over naive CXL. Although the benefit remains evident, the smaller margin arises from latency sensitivity rather than bandwidth limits. In this case, local DRAM cannot hold all latency-critical data, forcing some onto the slower AIC. This co-location increases access latency, which is alleviated when dual AICs are employed—providing additional bandwidth that effectively reduces latency. Section V-C presents detailed results for this case, where CXL-aware allocation can even surpass the DRAM-only baseline.
Finally, Figure 13(d) reports dual-GPU throughput for the 12 B model with one AIC. The naive CXL setup sustains 72%–91% of baseline performance, while CXL-aware allocation increases it to 82%–92%, offering up to 10% improvement over the naive policy. Nonetheless, concurrent bandwidth contention and latency sensitivity constrain further gains. The following section evaluates how the proposed allocation algorithm alleviates these bottlenecks under a dual-AIC environment.
V-C Performance Evaluation with Dual CXL AICs
The evaluation next turns to the dual-AIC scenario, designated as Config. B in Table II. Three configurations are used to establish the impact of CXL memory. The first is the baseline, where all data remains in local DRAM. The second is naive CXL, which pairs 128 GiB of DRAM with two 256 GiB AICs under a naive numactl --interleave=all policy. The third is CXL-aware allocation, which uses the same capacities but applies the proposed algorithm.
Figure 14(a) presents single-GPU throughput for a 7 B model across the same range of context lengths and batch sizes used earlier in the single-AIC configuration. While the naive CXL policy results in a 2% to 9% performance drop compared to the baseline, the proposed CXL-aware allocation restores performance to 100% of the baseline, showing that no performance is lost when CXL memory is managed intelligently. The algorithm achieves this by automatically interleaving latency-tolerant data across both AICs, thereby capitalizing on their aggregate bandwidth. These results demonstrate that with workload-aware allocation, a dual-card CXL configuration can fully match native DRAM performance.
Figure 14(b) shows dual-GPU throughput for a 7 B model across the same context lengths and batch sizes used in the single-AIC setup. The naive CXL policy lowers performance by 2% to 8% relative to the baseline, whereas the CXL-aware allocation restricts the loss to no more than 1%. In the single-AIC case shown in Figure 13(b), bandwidth contention remained a limiting factor even with CXL-aware placement. The dual-AIC setup alleviates this issue, allowing the algorithm to exploit the combined resources of both cards and largely eliminate the penalty, while the naive policy continues to falter due to poor placement.
Figure 14(c) illustrates single-GPU throughput for a 12 B model across the same range of context lengths and batch sizes. The naive CXL policy reduces throughput by 2% to 11% compared to the baseline, whereas the CXL-aware allocation restores performance to 100%–101% of the baseline. In the single-AIC case (Figure 13(c)), limited local DRAM capacity constrained performance even with CXL-aware placement. That limitation still exists in the dual-AIC setup, but the CXL-aware allocation mitigates its impact by leveraging the combined bandwidth of both cards alongside local DRAM. The latency-first greedy algorithm places latency-sensitive data in local DRAM whenever possible and assigns the remaining data across local DRAM and both AICs to aggregate bandwidth and minimize overall memory-access latency. The naive CXL policy, by contrast, continues to underperform due to unoptimized placement.
Finally, Figure 14(d) reports dual-GPU throughput for the 12 B model. The naive CXL policy reduces throughput by 2% to 9% relative to the baseline, while the CXL-aware allocation limits the loss to at most 1%. In the single-AIC case (Figure 13(d)), bandwidth contention and limited DRAM capacity remained bottlenecks even with CXL-aware placement. The dual-AIC configuration resolves these issues: CXL-aware allocation places latency-sensitive data appropriately and distributes remaining allocations across all CXL NUMA nodes to fully aggregate bandwidth. This automated interleaving enables both CPU and GPU workloads to exploit the combined bandwidth of the dual cards. The resulting performance, effectively matching the DRAM-only baseline, underscores the importance of this work to intelligently orchestrate underlying hardware resources.
@@ V-C Performance Evaluation with Dual CXL AICs
(c) 12B models in a single-GPU scenario
(d) 12B models in a dual-GPU scenario
Figure 15: Training latency decomposition for a long-context workload (32K context, batch size 4), showing our CXL-aware allocation consistently outperforms the naive CXL policy and nearly matches the DRAM-only baseline performance on both single- and dual-AIC systems. Detailed training latency comparison for five configurations, sorted by performance from right to left. The configurations are: (1) baseline, using local DRAM only, (2) naive CXL with single-AIC, (3) CXL-aware allocation with a single-AIC, (4) naive CXL with dual-AIC, and (5) CXL-aware allocation with dual-AIC.
V-D Latency Decomposition
While the previous section established end-to-end performance, this section breaks down the latency of FWD, BWD, and STEP for each configuration to further analyze latency and the impact of CXL-aware allocation. The system setup and configurations follow those used earlier: the single-AIC scenario corresponds to Config. A in Table II, and the dual-AIC scenario corresponds to Config. B. Five configurations are compared. The first is the baseline, where all data remain in local DRAM. The second is naive CXL (1 AIC), which combines 128 GiB of local DRAM with 512 GiB of CXL memory under a naive numactl --interleave=all policy. The third is CXL-aware allocation (1 AIC), which uses the same capacities but applies the proposed CXL-aware memory allocator. The fourth is naive CXL (2 AIC), pairing 128 GiB of DRAM with two 256 GiB AICs under the same interleave-all policy. The fifth is CXL-aware allocation (2 AIC), which employs the same capacities while applying the proposed extension and allocator.
Figure 15 presents a detailed latency breakdown of the training process into forward (FWD), backward (BWD), and optimizer-step (STEP) phases, revealing how memory configurations affect each stage. In the 7 B single-GPU scenario shown in Figure 15(a), FWD and BWD remain around $8.5s$ and $29.3s$ across all policies. The major separation arises in STEP: the baseline completes in $\approx 1.95s$, naive CXL (1 AIC) inflates it to $\approx 4.06s$, and naive CXL (2 AIC) reduces it slightly to $\approx 2.58s$. In contrast, CXL-aware allocation lowers STEP to $\approx 2.08s$ on one AIC and $\approx 1.96s$ on two AICs, essentially matching the baseline. This is because the naive interleave policy places latency-critical optimizer states on AIC memory, forcing frequent CPU accesses through a high-latency path. On the other hand, the CXL-aware allocation pins latency-sensitive tensors in local DRAM and pushes only bandwidth-tolerant data to CXL, eliminating this penalty.
Figure 15(b) shows the 7 B dual-GPU scenario, where bandwidth contention on the CXL interconnect becomes more pronounced. With a single AIC, both GPUs compete for bandwidth, slowing FWD and BWD even under CXL-aware allocation (1 AIC). This bottleneck highlights the limitation of a single CXL device in a multi-GPU environment. However, CXL-aware allocation (2 AIC) resolves this issue by intelligently distributing memory accesses across both AICs, providing sufficient aggregate bandwidth and restoring performance to within 0.2% of the DRAM-only baseline. For the larger 12 B model in a single-GPU setup (Figure 15(c)), FWD and BWD remain similar across configurations, but STEP dominates. The baseline STEP is $\approx 2.57s$; naive CXL (1 AIC) raises it to $\approx 6.11s$, while CXL-aware allocation (1 AIC) lowers it to $\approx 3.73s$. CXL-aware allocation (2 AIC) further cuts STEP to $\approx 2.59s$, effectively equal to the baseline. The small residual gap in the single-AIC aware case aligns with earlier findings: local DRAM cannot accommodate all latency-critical tensors for 12 B, so part of STEP still touches AIC memory. With only one card, this additional latency remains unavoidable.
Finally, the most demanding configuration, the 12 B model with dual GPUs, shown in Figure 15(d), demonstrates the combined effects of latency sensitivity and bandwidth contention. Single-AIC configurations suffer substantial degradation, with total latency increasing by up to 9% over the baseline because the AIC is fully saturated. Both FWD and BWD phases, as well as STEP, slow down. In contrast, CXL-aware allocation (2 AIC) effectively manages the hardware resources: by distributing allocations across local DRAM and both CXL NUMA nodes, it aggregates bandwidth while prioritizing latency-sensitive data, eliminating CXL-induced overhead and matching the performance of the DRAM-only baseline.
@@ VI Related Works
Offloading strategies have emerged as an effective means of breaking the GPU memory ceiling, enabling the training of models that would otherwise exceed on-device capacity by staging tensors in CPU DRAM or NVMe SSDs [16, 17, 18, 19, 20]. Among these, the ZeRO series [16, 17] has become the most widely adopted, owing to continuous maintenance that fixes bugs, adds support for new models, and preserves interoperability with complementary optimizations such as Flash-Attention [36] and Liger-Kernel [37]. Its reference implementation now underpins several popular training frameworks, including Accelerate [41] and MS-Swift [42]. Complementary work further broadens the design space of memory-centric training. Huang et al. [18] tailor an offloading mechanism specifically for large-scale language models, while Chen et al. [19] evaluate pragmatic orchestration policies that coordinate CPU and GPU memory during fine-tuning. Zeng et al. [20] extend the idea to fully heterogeneous environments, automatically balancing both compute and memory resources.
For CXL-attached memory, several tiered-memory systems (TMS) address the capacity challenge at the operating-system or hardware level. Representative examples include TPP [32], which classifies pages as hot or cold for placement, and NOMAD [43], which employs transactional page migration for asynchronous data movement. Other advanced systems, such as Colloid [44], dynamically balance traffic to equalize effective latency, while M5 [45] embeds hardware trackers in the CXL controller to provide fine-grained access statistics. Although these general-purpose TMS designs operate transparently without requiring application modifications, their workload-agnostic nature can lead to suboptimal performance for specialized applications such as LLM fine-tuning.
The analysis in this study identifies two weaknesses of applying a generic TMS to this workload. First, for data components transferred to the GPU, CXL-attached memory already offers bandwidth comparable to local DRAM. A TMS that migrates these pages to DRAM before a transfer would incur unnecessary page-movement overhead without improving performance. Second, the optimizer step exhibits a streaming access pattern in which every data element is updated once per iteration. This uniform pattern lacks the distinct hot or cold data regions that TMS architectures are designed to exploit. A generic tiering system would therefore be ineffective and could even introduce additional overhead through misguided migrations. In contrast, this study leverages application-specific knowledge to statically partition data, providing a more effective strategy for the predictable memory-access patterns of fine-tuning workloads.
@@ VII Conclusion
To address the latency challenge of CXL-attached memory relative to local DRAM during long-context LLM fine-tuning, particularly the performance degradation observed in CPU-intensive computations, this study identifies a fundamental limitation in existing deep learning frameworks: the lack of fine-grained control over memory allocation across heterogeneous memory systems. To overcome this limitation, two complementary approaches are proposed. First, a fine-grained memory allocation extension to PyTorch is developed to provide tensor-level control over system memory allocation policies, enabling fine-grained tensor management and facilitating future research with CXL-attached memory. Second, a CXL-aware memory allocator is introduced, which strategically places latency-sensitive data in local DRAM and latency-tolerant data in CXL-attached memory based on their access patterns. Together, these optimizations substantially mitigate the performance drawbacks of CXL-attached memory, consistently outperforming naive CXL adoption with up to 21% improvement across evaluated scenarios. In the dual-AIC configuration, the proposed method achieves near-baseline throughput, narrowing the gap with DRAM-only setups to within 1%, thereby demonstrating that CXL-attached memory can serve as a practical and high-performance solution for long-context LLM fine-tuning.